Vector v has a magnitude of 15 and a direction angle of 60 degrees. vector u has a magnitude of 25 and a direction angle of 140°.
What is the direction angle for the resultant vector v + u?

Answer by ikleyn(52780)

.

Vector v = (15*cos(60°),15*sin(60°)).


Vector u = (25*cos(140°),25*sin(140°)).


Vector u + v = (15*(cos(60°) + 25*cos(140°), 15*sin(60°) + 25*sin(140°))


If "a" is the direction angle for the resultant, then


    tan(a) = .


So, calculate tan(a) using this formula.


Then  a = arctan of this value.


Having clear instructions, you may complete calculations on your own.